# Particle transport; Dynamic Monte Carlo; Global variance reduction; Weight windows; Hybrid methods; Moment equations

### v1

@@ 1 Introduction
There are several ways to modify the particle weights and reduce variance in the solution. One option, implicit capture, decreases the weight of particles that would otherwise be absorbed, allowing them to survive through shielding. However, this can lead to many particles with extremely low weight, which will decrease the solution efficiency. Alternatively, splitting and rouletting can be used to keep particle weights constrained within some window about a chosen distribution with the goal of reducing variance.
@@
The choice of what function to use to define centers of weight windows depends on the type of problem being solved. One common choice is an adjoint scalar flux solution. This is useful in problems where the solution of interest is some functional, for instance, a detector response function. Using the adjoint solution redistributes computational particles toward the detector, decreasing variance at the location of the detector. The adjoint solution can be computed using a deterministic solver, leading to a hybrid method [1, 2].
@@
In global problems, the solution quality over the whole domain is of equal importance. In this class of problems, the adjoint solution method is not an efficient approach since an adjoint problem must be formulated and solved for each cell of a tally grid. Instead, it has been shown that constraining particle weights to a normalized forward flux solution results in reduced global variance in steady state neutron transport problems [3, 4]. This work used the hybrid low-order quasi-diffusion (LOQD) (aka variable Eddington factor) equations for the angular moments of the angular flux with the Eddington factor defining an exact closure. The Eddington factor is computed by Monte Carlo. Forward solution based weight windows have also been used in time dependent thermal radiative transfer (TRT). Weight windows defined using a solution of hybrid LOQD equations or the lagged Monte Carlo solution computed on the previous time step improved efficiency in these problems [5, 6].
Previously, the hybrid LOQD problem has been used to obtain the forward solution for both steady-state neutronics and time-dependent TRT. The second moment (SM) method is formulated with an alternative low-order problem for the scalar flux and current with exact linear closures [7]. The SM method has been used to develop hybrid implicit Monte Carlo (IMC) schemes for TRT [8, 9]. A hybrid SM method was shown to be effective with low particle counts per cell in steady-state transport problems [10, 11]. Additionally, the differential operator of the low-order second moment (LOSM) equations is self-adjoint which is advantageous when solving multi-dimensional problems.
@@
In this work, we formulate a new automatic weight windows technique for global variance reduction in time dependent particle transport problems. At each time step, the centers of weight windows are defined by the numerical solution of a hybrid Monte Carlo / deterministic problem based on the low-order SM (LOSM) equations. The hybrid LOSM equations are discretized in space by second-order finite volume scheme. The Crank-Nicolson time-integration method of the second-order accuracy is used for temporal approximation of the LOSM equations. The closures for the LOSM equations are computed by the Monte Carlo solution. One version of the developed method applied filtering techniques to the closures of the LOSM equations and their initial condition at each time step to mitigate the effect of stochastic noise. The moving average filter and Fourier filtering were both considered. We conducted a study of parameters which define the weight windows algorithm and used the results to motivate the choices in our Monte Carlo calculations. The performance of the proposed Monte Carlo algorithms with automatic weight windows are analyzed by means of an analytic transport benchmark. The developed weight windows algorithms have been implemented in the Monte Carlo / Dynamic Code (MC/DC) [12].
@@
The rest of the paper is organized as follows. Section 2 describes a general weight windows method for global time-dependent transport problems. Section 3 presents the weight window method for dynamic Monte Carlo calculations based on auxiliary solution of hybrid LOSM equations. The noise filtering techniques are described in Section 4. The numerical results are presented in Section 5. Section 6 concludes with a summary of the findings and directions for future work.
@@ 2 Weight Windows for Global Transport Problems
In non-analog Monte Carlo methods, particles are assigned a weight that can be changed to reduce variance. Weight windows are a way of controlling both the weights, and the spatial distribution of particles. In this scheme for global variance reduction, the weights are altered to obtain a more uniform particle distribution [3, 5]. The windows are applied when a particle undergoes an event or enters a new mesh cell. The main parameters of a weight window are its center, ceiling, and floor. We define the center of the weight window in the $i^{th}$ spatial cell as follows:
where $n$ is the index of the time step of dynamic Monte Carlo, $\tilde{\phi}_{i}^{n}$} is the approximate scalar flux in the $i^{th}$ cell at time $t^{n}$ evaluated by some numerical technique, $\varepsilon_{min}$ is the parameter which specifies the minimum center. The ceiling and floor of the weight window are given by
@@
Let consider a particle having a weight $w_{P}$. If the particle is in the $i^{th}$ cell during the $n^{th}$ time interval and $w_{P}>ww_{i}^{n,ceiling}$, the particle will be split into $N_{d}$ daughter particles, where $N_{d}~=~\lceil\frac{w_{P}}{ww_{i}^{n}}\rceil$. The original particles weight is distributed onto its daughter particles. If $w_{P}~<~ww_{i}^{n,floor}$, then the particle will be rouletted with survival probability $p=\frac{w_{P}}{ww_{i}^{n}}$ and if it survives, it will be assigned $w_{P}=ww_{i}^{n}$. Combined, these processes ensure that the weights of all particles in the cell fall within the window of that cell. Note that this process preserves the total weight of all particles and does not introduce any bias into the solution.
Since particles with weights higher than the ceiling are split, it is possible for cells with very low ceilings to experience excessive splitting. This also occurs in regions of the domain where the solution has a large gradient, such as wave fronts. When the window centers, $ww_{i}^{n}$, approach zero, the number of daughter particles, $\lceil\frac{w_{P}}{ww_{i}^{n}}\rceil$, grows by orders of magnitude. Time dependent weight windows can cause excessive splitting [6, 5]. In thermal radiative transfer, where waves are common feature, several modifications have been developed to artificially raise the window center near the wave front [5, 4]. The weight window center defined by Eq. (1) limits splitting by ensuring the window centers never fall below $\varepsilon_{min}$. The shape of the original window centers is preserved and scaled to fit above the minimum.
3 Hybrid Weight Window Method for Dynamic Monte Carlo Calculations
3.1 Hybrid Monte Carlo / Deterministic Method for Time-Dependent Transport Problem
@@
are cell-average values of the scalar flux and closure function ($i=1,\ldots,I$), respectively, and their values at domain boundaries are $\phi_{0}^{n}=\phi^{n}(x_{0})$, $F_{0}^{n}=F^{n}(x_{0})$, $\phi_{I+1}^{n}=\phi^{n}(x_{I+1})$, and $F_{I+1}^{n}~=~F^{n}(x_{I+1})$.
The closure terms of the discretized LOSM equations defined by $\{F_{i}^{n}\}_{i=0}^{I}$, $P_{L}^{n}$, and $P_{R}^{n}$ are computed by means of the Monte Carlo solution during the $n^{th}$ time step. Each time step is treated individually. The Monte Carlo solution is also used to calculate (i) the initial conditions for the $n^{th}$ time step defined by $\{\phi_{i}^{n-1}\}_{i=0}^{I+1}$ and $\{J_{i}^{n-1}\}_{i=0}^{I}$ and (ii) $\{F_{i}^{n-1}\}_{i=0}^{I+1}$ in the case of the CN scheme. This defines the hybrid Monte Carlo/ deterministic method based on the LOSM equations for global time-dependent transport calculations. We apply the numerical solution of the hybrid LOSM (HLOSM) equations (Eq. (13)) on each time step as an auxiliary solution $\{\tilde{\phi}_{i}^{n}\}_{i=1}^{I}$ to define centers of the weight windows according to Eq. (1).
@@
[…]
Figure 1: Tally options in Monte Carlo space-time cell
3.2 Updates of Weight Windows During Time Step
@@
During the $n^{th}$ time step ($t^{n-1}<t\leq t^{n}$), the weight windows are updated $u_{ww}^{n}$ times by means of the solution to the fully implicit HLOSM equations (Eq. (13)) using closures computed from Monte Carlo particle histories simulated before the update. Let $H^{n}$ be the total number of source particles on the $n^{th}$ time step. The $p^{th}$ update occurs after the number of histories reached the value given by
@@
where $f_{p}$ is the fraction of total number particle histories specified for the $p^{th}$ update.
@@ 4 Statistical Noise Filtering of Closures and Initial Conditions
where $\tilde{\chi}_{j}$ is the filtered $j^{th}$ element of the grid function and $k\in\mathbb{N}$ is the parameter of the filter base which specifies the spatial interval for averaging. A filter with large interval results in more smoothing at the expense of preserving details in the grid function. The resolution can be adjusted by optimizing the filter base. Near boundaries of the spatial domain, the filter base extends outside of the domain. The parameter $k$ is adaptively decreased to avoid this issue.
@@ 5.1 Test Problem
To analyze the proposed hybrid weight window (HWW) method, we use a dynamic neutron transport benchmark problem with a semi-analytic solution [15]. This is a 1D slab geometry problem for infinite homogeneous medium. The material cross sections are $\Sigma_{s}=\frac{1}{3}\,cm^{-1}$, $\Sigma_{f}~=~\frac{1}{3}\,cm^{-1}$, $\Sigma_{t}=1\,cm^{-1}$, and $\nu_{f}=2.3$. The speed of particles is $v=1\frac{cm}{s}$. A point source is located in the domain center at $x=0$ and produces a particle impulse at $t=0$ with strength 1 $\frac{particle}{cm^{3}\,s}$. There are no particles in the domain at $t=0$. The time interval of the problem is $0\leq t\leq 20s$. The spatial mesh consists of 201 uniform cells covering the region $20.5cm\leq x\leq 20.5cm$. The test is calculated with the time step $\Delta t=1$ s.
@@ 5.2 Analysis of Weight Window Parameters
The HWW method has several parameters. First, we consider the parameters $u_{ww}^{n}$ which define the number of updates of weight window centers using recalculated solution of the HLOSM equations. Determining when and how many times to do these updates is based on two competing interests. Updating too soon means that the updated closures will have larger statistical error. On the other hand, if the weight windows are updated too late, the update could have little impact on the final tallies. To demonstrate effects of weight windows updates on the Monte Carlo solution, we present the results for the test problem computed with different numbers of updates $u_{ww}^{n}$. The test is solved using $H^{n}=10^{4}$ particle histories on each time step and with $f_{p}=\frac{p}{u_{ww}^{n}+1}$. Note that the set $\{f_{p}\}_{p=1}^{u_{ww}^{n}}$ defines frequency of updates in terms of number of particle histories. The parameters of weight windows are $\rho=1.25$ and $\epsilon_{min}=10^{-3}$. The center of weight windows are computed by means of numerical solution of the HLOSM equations discretized with the CN temporal scheme and without filtering of closures and initial conditions. Figure 4 presents the relative $L_{2}$-norm of the error in the scalar flux computed by Monte Carlo algorithm, $\langle\phi\rangle^{n}$, over two different time steps ($n=4,19$). The relative error is calculated using the benchmark solution, $\phi_{ref}$. The results in Figure 4 show that the error decrease as $\frac{1}{\sqrt{H_{p}^{n}}}$ in the case of a small number of updates executed with relatively large fractions of particle histories. We also present the results of calculations with a very large number of updated $u_{ww}=8,16$. Such a number of updates is not efficient, but these cases were used to analyze feedback effects of frequent updates of weight windows on Monte Carlo solution. The results show that when the number of updates is large, the Monte Carlo solution converges but there are stages over which the error increases. This indicates potential instabilities of computations with large number and frequent updates of weight windows. Figure 5 shows the relative error in the solution of HLOSM equations, $\phi_{hlosm}$, which is used as an auxiliary solution $\tilde{\phi}$ to define weight windows (Eq. (1)) in Monte Carlo calculations presented in Figure 4. As more histories run, the error in closures decreases. This leads to improvement in accuracy of numerical solution of the HLOSM equations. At some number of histories, the stochastic component of error in HLOSM solution becomes smaller then the error due to discretization in space and time. As a result, the error limits to some value. The point where the stochastic error of the hybrid solution becomes smaller than the effects of discretization error depends on the mesh and number of histories [10].
@@
The next parameter of interest is $\varepsilon_{min}$ which defines the minimum value of the weight window center to limit the amount of splitting. Decreasing the parameter $\epsilon_{min}$ leads to better resolution of the wave front at the expense of additional particles and therefore increases computational cost. Figure 6 shows the spatial distribution of particle tracks per source particle at time steps $n=10,20$ solved with $H^{n}=10^{4}$ source particles and window windows defined with $\rho=1.25$ and various values of $\epsilon_{min}$. The analog particle tracks in Monte Carlo simulations without weight windows technique are peaked in the center, near the initial source of particles. This leads to underpredicting of wavefront position. Applying weight windows improves the resolution of the wave front with Monte Carlo particles. At the time step n=20, the best resolution is achieved with $\epsilon_{min}=10^{-6}$. In this case, the total particle population at the end the time step grows by orders of magnitude. One source particle with initial weight $w=1$ would need to become approximately $10^{6}$ particles each with weight $w\approx 10^{-6}$.

### v2

@@ 1 Introduction
There are several ways to modify the particle weights and reduce variance in the solution. One option, implicit capture, decreases the weight of particles that would otherwise be absorbed, allowing them to survive through shielding. However, this can lead to many particles with extremely low weight, which will decrease the solution efficiency. Alternatively, splitting and rouletting can be used to keep particle weights constrained within some window about a chosen distribution with the goal of reducing variance [1, 2].
@@
The choice of what function to use to define centers of weight windows depends on the type of problem being solved. For problems where a detector response functional is desired, the optimal function is the adjoint flux solution. Using the adjoint solution redistributes computational particles toward the detector, decreasing variance at the location of the detector. The adjoint solution can be computed using a deterministic solver, leading to a hybrid method [3, 4].
@@
In global problems, the solution quality over the whole domain is of equal importance. In this class of problems, the adjoint approach requires two deterministic solves, one forward solve that is used to define the adjoint source and one adjoint solve to obtain the importance function [5]. Using a normalized forward flux solution to define weight windows has been shown to reduce global variance in steady state neutron transport problems [6, 7]. This work used the hybrid low-order quasi-diffusion (LOQD) (aka variable Eddington factor (VEF)) equations for the angular moments of the angular flux with the Eddington factor defining an exact closure [8, 9]. The Eddington factor is computed by Monte Carlo.
In time dependent thermal radiative transfer (TRT) problems, the forward flux LOQD solution has been used to define hybrid weight windows (HWW) [10]. For these HWW, the LOQD (VEF) equations were discretized in time with a scheme of the first order of accuracy. Multiphysics methods that are second order in time have also been developed for TRT problems [11]. The Second Moment (SM) method has been used to develop hybrid implicit Monte Carlo (IMC) schemes for TRT [12, 13]. The low-order SM (LOSM) equations are formulated for the scalar flux and current with exact linear closures [14].
@@
In this paper, we present a new automatic HWW technique for global variance reduction in time dependent particle transport problems. At each time step, the centers of weight windows are defined by the numerical solution of a hybrid Monte Carlo / deterministic problem based on the LOSM equations. A hybrid SM method was shown to be effective with low particle counts per cell in steady-state transport problems [15, 16]. Additionally, the differential operator of the low-order second moment (LOSM) equations is self-adjoint, which is advantageous when solving multi-dimensional problems. The hybrid LOSM equations are discretized in space by second-order finite volume scheme. The Crank-Nicolson time-integration method of the second-order accuracy is used for temporal approximation of the LOSM equations. The closures for the LOSM equations are computed by the Monte Carlo solution. One version of the developed method applied filtering techniques to the closures of the LOSM equations and their initial condition at each time step to mitigate the effect of stochastic noise. The moving average filter and Fourier filtering were both considered. We conducted a study of parameters which define the weight windows algorithm and used the results to motivate the choices in our Monte Carlo calculations. The performance of the proposed Monte Carlo algorithms with automatic weight windows are analyzed by means of an analytic transport benchmark. The developed weight windows algorithms have been implemented in the Monte Carlo / Dynamic Code (MC/DC) [17].
@@
The rest of the paper is organized as follows. Section 2 describes a general weight windows method for global time-dependent transport problems. Section 3 presents the weight window method for dynamic Monte Carlo calculations based on auxiliary solution of hybrid LOSM equations. The noise filtering techniques are described in Section 4. In Section 5 are numerical results for the parameter study as well as an analysis of the performance of HWW. Section 6 concludes with a summary of the findings and directions for future work.
@@ 2 Weight Windows for Global Transport Problems
In non-analog Monte Carlo methods, particles are assigned a weight that can be changed to reduce variance. Weight windows are a way of controlling both the weights, and the spatial distribution of particles. In this scheme for global variance reduction, the weights are altered to obtain a more uniform particle distribution [6, 10]. The windows are applied when a particle undergoes an event or enters a new mesh cell. The main parameters of a weight window are its center, ceiling, and floor. The center of the weight window in the $i^{th}$ spatial cell is usually defined as follows [6]:
$ww_{i}^{n}=\frac{{\tilde{\phi}_{i}^{n}}}{\max\limits_{i}\tilde{\phi}_{i}^{n}},$ (1)
where $n$ is the index of the time step of dynamic Monte Carlo, $\tilde{\phi}_{i}^{n}$} is the approximate scalar flux in the $i^{th}$ cell at time $t^{n}$ evaluated by some numerical technique. The ceiling and floor of the weight window are given by
@@
Let us consider a particle having a weight $w_{P}$. If the particle is in the $i^{th}$ cell during the $n^{th}$ time interval and $w_{P}>ww_{i}^{n,ceiling}$, the particle will be split into $N_{d}$ daughter particles, where $N_{d}=\lceil\frac{w_{P}}{ww_{i}^{n}}\rceil$. The original particles weight is distributed onto its daughter particles. If $w_{P}<ww_{i}^{n,floor}$, then the particle will be rouletted with survival probability $p=\frac{w_{P}}{ww_{i}^{n}}$ and if it survives, it will be assigned $w_{P}=ww_{i}^{n}$. Combined, these processes ensure that the weights of all particles in the cell fall within the window of that cell. Note that this process preserves the total weight of all particles and does not introduce any bias into the solution.
Since particles with weights higher than the ceiling are split, it is possible for cells with very low ceilings to experience excessive splitting. This also occurs in regions of the domain where the solution has a large gradient, such as wave fronts. When the window centers, $ww_{i}^{n}$, approach zero, the number of daughter particles, $\lceil\frac{w_{P}}{ww_{i}^{n}}\rceil$, grows by orders of magnitude. Time dependent weight windows can cause excessive splitting [18, 10]. In thermal radiative transfer, where waves are common feature, several modifications have been developed to artificially raise the window center near the wave front [10, 7]. To limit excessive splitting in this work, we define a modified window center given by
where $\varepsilon_{min}$ is a parameter that sets the minimum value of the window centers. Using this modification, the shape of the original window centers is preserved and scaled to fit above the minimum.
3 Hybrid Monte Carlo / Deterministic Method for Time-Dependent Transport Problem
@@
are cell-average values of the scalar flux and closure function ($i=1,\ldots,I$), respectively, and their values at domain boundaries are $\phi_{0}^{n}=\phi^{n}(x_{0})$, $F_{0}^{n}=F^{n}(x_{0})$, $\phi_{I+1}^{n}=\phi^{n}(x_{I+1})$, and $F_{I+1}^{n}=F^{n}(x_{I+1})$.
The closure terms of the discretized LOSM equations defined by $\{F_{i}^{n}\}_{i=0}^{I}$, $P_{L}^{n}$, and $P_{R}^{n}$ are computed by means of the Monte Carlo solution during the $n^{th}$ time step. Each time step is treated individually. The Monte Carlo solution is also used to calculate (i) the initial conditions for the $n^{th}$ time step defined by $\{\phi_{i}^{n-1}\}_{i=0}^{I+1}$ and $\{J_{i}^{n-1}\}_{i=0}^{I}$ and (ii) $\{F_{i}^{n-1}\}_{i=0}^{I+1}$ in the case of the CN scheme. This defines the hybrid Monte Carlo/ deterministic method based on the LOSM equations for global time-dependent transport calculations. We apply the numerical solution of the hybrid LOSM (HLOSM) equations (Eqs. (14)) on each time step as an auxiliary solution $\{\tilde{\phi}_{i}^{n}\}_{i=1}^{I}$ to define centers of the weight windows according to Eq. (3).
@@
[…]
Figure 1: Tally options in Monte Carlo space-time cell
@@
During the $n^{th}$ time step ($t^{n-1}<t\leq t^{n}$), the weight windows are updated $u_{ww}^{n}$ times by means of the solution to the fully implicit HLOSM equations (Eqs. (14)) using closures computed from Monte Carlo particle histories simulated before the update. Let $H^{n}$ be the total number of source particles on the $n^{th}$ time step. The $p^{th}$ update occurs after the number of histories reached the value given by
@@
where $f_{p}$ is the fraction of the total number of particle histories specified for the $p^{th}$ update.
@@ 4 Statistical Noise Filtering of Closures and Initial Conditions
where $\tilde{\chi}_{j}$ is the filtered $j^{th}$ element of the grid function and $k\in\mathbb{N}$ is the parameter of the filter base which specifies the spatial interval for averaging. A filter with a large interval results in more smoothing at the expense of preserving details in the grid function. The resolution can be adjusted by optimizing the filter base. Near boundaries of the spatial domain, the filter base extends outside of the domain. The parameter $k$ is adaptively decreased to avoid this issue.
@@ 5 Numerical Results
In this section, we present numerical results for various aspects of the hybrid problem. First, a description of the test is given. Hybrid solutions computed with CN and BE time discretizations are shown, illustrating the advantage of the higher order CN discretization. The results of the parameter study in $u_{ww}^{n}$, $\epsilon_{min}$, and $\rho$ are then presented, motivating the choice of parameters. Following that is an analysis of the quality of the hybrid LOSM solution, and comparison to the hybrid Monte Carlo solution. The effects of filtering techniques on the closure are also presented here. Finally, a comparison in terms of particle track density, relative variance, and figure of merit between analog Monte Carlo, and hybrid weight windows is presented.
@@ 5.1 Test Problem
To analyze the proposed hybrid weight window method, we use a dynamic neutron transport benchmark problem with a semi-analytic solution [21]. This is a 1D slab geometry problem for infinite homogeneous medium. The material cross sections are $\Sigma_{s}=\frac{1}{3}\,cm^{-1}$, $\Sigma_{f}=\frac{1}{3}\,cm^{-1}$, $\Sigma_{t}~=~1~cm^{-1}$, and $\nu_{f}=2.3$. The speed of particles is $v=1\frac{cm}{s}$. A point source is located in the domain center at $x=0$ and produces a particle impulse at $t=0$ with strength 1 $\frac{particle}{cm^{3}\,s}$. There are no particles in the domain at $t=0$. The time interval of the problem is $0\leq t\leq 20s$. The spatial mesh consists of 201 uniform cells covering the region $20.5cm\leq x\leq 20.5cm$. The test is calculated with the time step $\Delta t=1$ s.
@@ 5.3 Analysis of Weight Window Parameters
The hybrid weight window method has several parameters. First, we consider the parameters $u_{ww}^{n}$ which define the number of updates of weight window centers using recalculated solution of the HLOSM equations. Determining when and how many times to do these updates is based on two competing interests. Updating too soon means that the updated closures will have larger statistical error. On the other hand, if the weight windows are updated too late, the update could have little impact on the final tallies. To demonstrate effects of weight windows updates on the Monte Carlo solution, we present the results for the test problem computed with different numbers of updates $u_{ww}^{n}$. The test is solved using $H^{n}=10^{4}$ particle histories on each time step and with $f_{p}=\frac{p}{u_{ww}^{n}+1}$. Note that the set $\{f_{p}\}_{p=1}^{u_{ww}^{n}}$ defines frequency of updates in terms of number of particle histories. The parameters of weight windows are $\rho=1.25$ and $\epsilon_{min}=10^{-3}$. The center of weight windows are computed by means of numerical solution of the HLOSM equations discretized with the CN temporal scheme and without filtering of closures and initial conditions. Figure 4 presents the relative $L_{2}$-norm of the error in the scalar flux computed by Monte Carlo algorithm, $\langle\phi\rangle^{n}$, over two different time steps ($n=4,19$). The relative error is calculated using the benchmark solution, $\phi_{ref}$. The results in Figure 4 show that the error decreases as $\frac{1}{\sqrt{H_{p}^{n}}}$ in the case of a small number of updates executed with relatively large fractions of particle histories. We also present the results of calculations with a very large number of updated $u_{ww}=8,16$. Such a number of updates is not efficient, but these cases were used to analyze feedback effects of frequent updates of weight windows on Monte Carlo solution. The results show that when the number of updates is large, the Monte Carlo solution converges, but there are stages over which the error increases. This indicates potential instabilities of computations with large number and frequent updates of weight windows. Figure 5 shows the relative error in the solution of HLOSM equations, $\phi_{hlosm}$, which is used as an auxiliary solution $\tilde{\phi}$ to define weight windows (Eq. (3)) in Monte Carlo calculations presented in Figure 4. As more histories run, the error in closures decreases. This leads to improvement in accuracy of numerical solution of the HLOSM equations. At some number of histories, the stochastic component of error in HLOSM solution becomes smaller then the error due to discretization in space and time. As a result, the error limits to some value. The point where the stochastic error of the hybrid solution becomes smaller than the effects of discretization error depends on the mesh and number of histories [15].
The next parameter of interest is $\varepsilon_{min}$ which defines the minimum value of the weight window center to limit the amount of splitting. Decreasing the parameter $\epsilon_{min}$ leads to better resolution of the wave front at the expense of additional particles and therefore increases computational cost. Figure 6 shows the spatial distribution of particle tracks per source particle at time steps $n=10,20$ solved with $H^{n}=10^{4}$ source particles and window windows defined with $\rho=1.25$ and various values of $\epsilon_{min}$. The analog particle tracks in Monte Carlo simulations without weight windows technique are peaked in the center, near the initial source of particles. This leads to underpredicting of wavefront position. Applying weight windows improves the resolution of the wave front with Monte Carlo particles. At the time step n=20, the best resolution is achieved with $\epsilon_{min}=10^{-6}$. In this case, the total particle population at the end of the time step grows by orders of magnitude. One source particle with initial weight $w=1$ would need to become approximately $10^{6}$ particles each with weight $w\approx 10^{-6}$.